OpenStudy (anonymous):

find the y-intercepts of f(x)=3e^(-.5x) -9. I really need an explanation on this one

OpenStudy (agent47):

y-intercept means x=0, so plug in 0 instead of x and solve for y.

OpenStudy (agent47):

f(0)=3*e^(0)-9=3*1-9=-6

OpenStudy (agent47):

point is: (0;-6)

OpenStudy (anonymous):

oh ok thanks. that makes sense now